OpenStudy (anonymous):

find the equation for the line passing through (3,-8) and (1,-2)

OpenStudy (anonymous):

1)Find slope 2)put into point-slope form 3)change to slope-intercept form

OpenStudy (anonymous):

1)Find Slope: y - y over x - x -8 -(-2) = -6 3 - 1 = 2 -6/2 = -3/1 slope: -3

OpenStudy (anonymous):

2) Point slope: y - y1 = m(x - x1) y - (-8) = -3(x - 3) y + 8 = -3x + 9

OpenStudy (anonymous):

3) Change to Slope-intercept form y + 8 = -3x + 9 -8 y = -3x + 9 - 8 y = -3x + 1 Slope-intercept form: y = -3x + 1

OpenStudy (anonymous):

Hope I helped!!

OpenStudy (anonymous):

thanks u helped alot